A solution containing 30% chicken broth is to be
mixed with a solution containing 50% chicken broth
to make 200L containing 42% chicken broth. How
much of each solution should be mixed

Answer by stanbon(75887)

A solution containing 30% chicken broth is to be
mixed with a solution containing 50% chicken broth
to make 200L containing 42% chicken broth. How
much of each solution should be mixed
-------------
Equation:
0.30x + 0.50(200-x) = 0.42*200
-----------
30x + 50*200 - 50x = 42*200
-20x = -8*200
x = 80 L (amt. of 30% solution needed)
----
200-x = 120 L (amt. of 50% solution needed)
==============================================
Cheers,
Stan H.